# The Gauge Theory Bootstrap: Predicting pion dynamics from QCD

### v2

@@ 1 Motivation and summary
In [1, 2, 3] we proposed the Gauge Theory Bootstrap (GTB), a method to understand this region from a purely theoretical point of view. The idea is to write the most general S-matrix, form factors and current spectral densities that satisfy basic physical properties and add more and more constraints based on matching of the physics to the known UV and IR. In that way we expect to converge to a unique solution parameterized by the UV and IR properties. In this paper we get much closer to this main goal by adding further IR constraints on the spectral density that greatly reduce the space of solutions and, further, by creating an iterative procedure that pushes the constraints towards saturation and converges to a unique solution. Instead of using a functional to map out the space of solutions, and extract extremal amplitudes as we did in previous work, we can remove the functional and allow the program to choose an arbitrary feasible point and then use this iterative procedure to converge to a solution. This differs from the standard S-matrix bootstrap procedure that uses kinematic constraints to put bounds on theory parameters, instead we introduce dynamical information [^1] We are grateful to A. Vainshtein for emphasizing this point. from the UV and IR to compute the S-matrix. In that way, we perhaps go back to the original bootstrap philosophy of the 1960’s [13, 14] but augmented with the microscopic dynamical information that was not known at that time. In that work [13, 14] the dynamical information was the idea of self-consistency, or hadron democracy: for example the $\rho$ meson gave rise to an attractive potential between pions which in turn gave rise to a resonance, the $\rho$ meson itself (see for example [15, 16]). Although interesting, such ideas did not match the experimental results and were superseded by QCD. In the recent bootstrap revival [17, 18, 19, 20], powerful numerical methods were used to put bounds on couplings and other parameters of generic quantum field theories, including the effective field theory of pions [21, 22, 23, 24, 25, 26, 27, 28, 29]. This also extended to form factors and spectral densities (mostly in 2d theories) [30, 31, 32, 33, 34]. However it is clear that general kinematical constraints, albeit non-trivial, still allow an enormous number of theories and that dynamical information, here provided by QCD, is necessary to identify the correct theory. [^2] In [32], dynamical information about the $\phi^{4}$ theory was extracted using Hamiltonian truncation and used as input in bootstrap; in [33], the central charge of $2d$ Ising CFT is used as an input. In that way, without using experimental input, we can predict and compute low energy constants, partial waves, form factors, resonance spectrum, etc. For example in the recent work [35] this method was used as a tool to study the g2g-2 factor contribution from hadrons. Notice also that the only assumption on the spectrum is that there are no bound states. The number of resonances together with their masses and widths are results of the computation based on the UV input from pQCD and matching with low energy pions.
@@ 2.1 Low energy spectral density
In appendix B we describe a parameterization of the spectral densities that takes into account this behavior. This is the first addition to the procedure and plays a crucial role because the spectral density is now also constrained in the IR and not only in the UV (by the SVZ sum rules).
@@ 2.2 Iterative “Watsonian” procedure
In (2.7) below, where only two pion states are considered, this saturation should certainly be exact below the four pion threshold and approximately above. In general, we adopt the philosphy of always look for saturation of the constraints and, if that is not possible to further add pion states and QCD operators to the matrix. Once a matrix such as (2.7) is chosen, to achieve the desired saturation, we now introduce an iterative procedure that leads to a particular solution to the GTB constraints.
Iterative procedures in the S-matrix bootstrap were introduced initially in [42] to get closer to the boundary of the allowed space. The main idea is that, denoting the bootstrap variables as $x_{j}$, and given a constraint $\Phi_{A}(x_{j})\leq 0$, after finding a solution $x_{j}=\bar{x}_{j}$, we can improve the solution by maximizing again in the direction of the gradient of the constraint. The unitarity constraint reads
@@
where $h^{I}_{\ell}(s)$ is the normalized partial wave where
with $\delta_{\ell}^{I}$ the usual phase shift. It is known, qualitatively, that pion scattering tends to saturate unitarity even at energies larger than the four pion threshold.666at higher energy we could incorporate some method of particle production such as was done in [43, 44] or via resonance decay as suggested in [3] If we want to approach further the unitarity saturation region, the new functional to maximize, for the next iteration, is
@@
where old are the values from the previous iteration, and new are the values in the maximization that is being performed. In the case of two dimensions it was very successful in converging to the desired solutions. One might expect that the same happens here. And although this is true, once we include form factors and spectral densities in the bootstrap, Watson theorem provides a new natural iterative functional for our purpose. Watson theorem establishes that, when unitarity is saturated, the phase of the form factors agrees with the phase shift of the corresponding partial wave. The procedure then replaces the phase of the previous (old) partial wave by the phase of the previous (old) form factor when computing the functional, i.e., in (2.5) we replace:
@@
Namely instead of just using the old partial wave to determine the functional, we use its modulus and replace its phase by the phase of the form factor. Such iteration will tend to converge to unitarity saturation while matching the phase of the form factor with the phase shift. From a physical point of view, the UV information incorporated into the form factors and spectral densities is conveyed to the phase shifts and partial waves. At the same time the crossing and unitarity constraints on the partial waves feed back to the form factors. More precisely, following the form factor bootstrap [30, 31] we introduce a $3\times 3$ matrix of scalar products:
@@
$\begin{array}[]{c c}&\begin{array}[]{c c c}\mbox{$|\mbox{in}\rangle$}_{P,\ell}&\mbox{$|\mbox{out}\rangle$}_{P,\ell}&{\cal O}_{\ell,P}\mbox{$|0\rangle$}\\ \end{array}\\ \begin{array}[]{c c c}\mbox{$\langle\mbox{in}|$}_{P^{\prime},\ell}\\ \mbox{$\langle\mbox{out}|$}_{P^{\prime},\ell}\\ \mbox{$\langle 0|$}{\cal O}^{\dagger}_{P^{\prime},\ell}\end{array}&\left(\begin{array}[]{c c c}1&\ S_{\ell}(s)&\ \ {\cal F}_{\ell}\\ S_{\ell}^{*}(s)&\ 1&\ \ {\cal F}^{*}_{\ell}\\ {\cal F}^{*}_{\ell}&\ {\cal F}_{\ell}&\ \ \rho_{\ell}(s)\end{array}\right)\end{array}\succeq 0,\;\;\forall I,\ell,\ s>4m^{2}_{\pi},\;\;$ (2.7)
@@
where we removed an overall factor $\delta^{4}(P-P^{\prime})$, namely the identity in the center of mass variable. The in- and out- states are two pion states and ${\cal O}_{\ell,P}$ is an integrated current. The elements of the matrix are the scattering matrix $S_{\ell}^{I}$, the form factors ${\cal F}_{\ell}^{I}$ and the spectral density $\rho_{\ell}^{I}$. When unitarity is saturated and the phase shift is equal to the phase of the form factor (Watson’s theorem) this matrix has two zero eigenvalues. That means that two linear combinations of the states vanish. One simply means that the in and out states are the same (up to a phase) if there is no particle production. The other, interestingly, means that we have a way to write ${\cal O}_{\ell,P}\mbox{$|0\rangle$}$ as a two pion state. So, from a physical perspective, one can understand the Watsonian iteration as a way to write pQCD states (currents acting on vacuum) in terms of pion states and vice versa.
@@ 2.3 Asymptotic form factors
This new improved procedure highlights the need for more precise high energy inputs. This is not a problem for the SVZ sum rules [38, 39, 40] but it is a problem for the form factors. The theoretical computation of form factors at high energy is well established [45] and proceeds similarly to the better known case of Deep Inelastic Scattering. However, there is less validation from experiments specially in the case of the pion. Qualitatively, at high energy, the pion is highly boosted but in the transverse directions it has the same size given by its radius $R_{\pi}\sim\frac{1}{f_{\pi}}$. The probability of a probing particle hitting a quark is inversely proportional to the transverse area. That is the reason that $f_{\pi}$ appears in the asymptotic formulas. More technically, the parton distribution functions are required, a low energy information. At high energy the results of Brodsky-Lepage [36, 46, 47, 3] allows us to determine form factors as
@@ 6 Thermodynamics of a dilute gas of pions, second virial coefficients and density of pion pairs
A quantity that is of interest for pions produced in high energy nuclear collisions [^9] We are grateful to Fuqiang Wang for a discussion on this topic. [66], when the quark gluon plasma cools down, is the difference in the pairs of pions with opposite (electric) charge $N_{\mbox{os}}$ minus the number of pairs with the same sign $N_{\mbox{ss}}$. For free pions this difference vanishes if we compute it as $N_{\mbox{os}}-N_{\mbox{ss}}=2\langle\langle n_{+}\rangle\rangle\langle\langle n_{-}\rangle\rangle-\langle\langle n_{+}\rangle\rangle\langle\langle n_{+}\rangle\rangle-\langle\langle n_{-}\rangle\rangle\langle\langle n_{-}\rangle\rangle$ since $\langle\langle n_{+}\rangle\rangle=\langle\langle n_{-}\rangle\rangle$ The interactions produce a difference in the number of pairs giving, at the lowest order of the virial expansion:
$N_{\mbox{os}}-N_{\mbox{ss}}$ $=$ $V\,z^{2}\int_{2m_{\pi}}^{\infty}n_{os-ss}(M)dM$ (6.19)
$n_{os-ss}(M)$ $=$ $\frac{M^{2}}{2\pi^{2}\beta}\,K_{2}(\beta M)\sum_{\ell}{}^{{}^{\prime}}\frac{2\ell+1}{\pi}\frac{\partial}{\partial M}\left\{\frac{1}{3}\delta^{I=0}_{\ell}-\frac{5}{6}\delta^{I=2}_{\ell}+\frac{1}{2}\delta^{I=1}_{\ell}\right\}$
where the $\sum^{\prime}$ indicates that we only sum (for each term) the angular momentum allowed for that isospin (even for $I=0,2$ and odd for $I=1$). Also, $M$ is the invariant mass of the pair ($M^{2}=(p_{1}+p_{2})^{2}$), and, for the case of zero chemical potential ($\mu=0)$, we have $z=1$. The results are depicted in fig. 15. The $\rho$ resonance peak coming from $\frac{\partial\delta_{1}^{1}}{\partial M}$ is evident. It can be interpreted as an increase in the density of states near the $\rho$ mass or as $\rho$ mesons being created and destroyed leading in both cases to an extra density of pairs of opposite sign.

### v3

@@ 1 Motivation and summary
In [1, 2, 3] we proposed the Gauge Theory Bootstrap (GTB), a method to understand this region from a purely theoretical point of view. The idea is to write the most general S-matrix, form factors and current spectral densities that satisfy basic physical properties and add more and more constraints based on matching of the physics to the known UV and IR. In that way we expect to converge to a unique solution parameterized by the UV and IR properties. In this paper we get much closer to this main goal by adding further IR constraints on the spectral density that greatly reduce the space of solutions and, further, by creating an iterative procedure that pushes the constraints towards saturation and converges to a unique solution. Instead of using a functional to map out the space of solutions, and extract extremal amplitudes as we did in previous work, we can remove the functional and allow the program to choose an arbitrary feasible point and then use this iterative procedure to converge to a solution. This differs from the standard S-matrix bootstrap procedure that uses kinematic constraints to put bounds on theory parameters, instead we introduce dynamical information [^1] We are grateful to A. Vainshtein for emphasizing this point. from the UV and IR to compute the S-matrix. In that way, we perhaps go back to the original bootstrap philosophy of the 1960’s [13, 14] but augmented with the microscopic dynamical information that was not known at that time. In that work [13, 14] the dynamical information was the idea of self-consistency, or hadron democracy: for example the $\rho$ meson gave rise to an attractive potential between pions which in turn gave rise to a resonance, the $\rho$ meson itself (see for example [15, 16]). Although interesting, such ideas did not match the experimental results and were superseded by QCD. In the recent bootstrap revival [17, 18, 19, 20], powerful numerical methods were used to put bounds on couplings and other parameters of generic quantum field theories, including the effective field theory of pions [21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. This also extended to form factors and spectral densities (mostly in 2d theories) [31, 32, 33, 34, 35]. However it is clear that general kinematical constraints, albeit non-trivial, still allow an enormous number of theories and that dynamical information, here provided by QCD, is necessary to identify the correct theory. [^2] In [33], dynamical information about the $\phi^{4}$ theory was extracted using Hamiltonian truncation and used as input in bootstrap; in [34], the central charge of $2d$ Ising CFT is used as an input. In that way, without using experimental input, we can predict and compute low energy constants, partial waves, form factors, resonance spectrum, etc. For example in the recent work [36] this method was used as a tool to study the g2g-2 factor contribution from hadrons. Notice also that the only assumption on the spectrum is that there are no bound states. The number of resonances together with their masses and widths are results of the computation based on the UV input from pQCD and matching with low energy pions.
@@ 2.1 Low energy spectral density
In appendix B we describe a parameterization of the spectral densities that takes into account this behavior. This is the first addition to the procedure and plays a crucial role because the spectral density is now also constrained in the IR and not only in the UV (by the SVZ sum rules as explained in section 2.4 and appendix C).
@@ 2.2 Iterative “Watsonian” procedure
In (2.9) below, where only two pion states are considered, this saturation should certainly be exact below the four pion threshold and approximate above. In general, we adopt the philosphy of always look for saturation of the constraints (i.e. matrix with zero modes) and, if that is not possible to further add pion states and QCD operators to the matrix. Once a matrix such as (2.9) is chosen, to achieve the desired saturation, we now introduce an iterative procedure that leads to a particular solution to the GTB constraints.
Iterative procedures in the S-matrix bootstrap were introduced initially in [43] to get closer to the boundary of the allowed space. The main idea is that, denoting the bootstrap variables as $x_{j}$, and given a constraint $\Phi_{A}(x_{j})\leq 0$, after finding a solution $x_{j}=\bar{x}_{j}$, we can improve the solution to saturate the constraint by maximizing again in the direction of the gradient of the constraint, namely a functional linear in the variables $x_{j}$:
$\mathcal{F}=\sum_{A,j}\left.\frac{\partial\Phi_{A}}{\partial x_{j}}\right|_{\bar{x}_{j}}x_{j}$ (2.3)
In terms of the normalized partial wave $h^{I}_{\ell}(s)$ where
with $\delta_{\ell}^{I}$ the usual phase shift, the unitarity constraint reads
@@
It is known, qualitatively, that 2-to-2 pion scattering tends to saturate unitarity even at energies larger than the four pion threshold.666at higher energy we could incorporate some method of particle production such as was done in [44, 45] or via resonance decay as suggested in [3] If we want to approach further the unitarity saturation, we can linearize the constraint (2.5) and the new functional to maximize, for the next iteration, is
@@
where old are the values from the previous iteration, and new are the values in the maximization that is being performed. In the case of two dimensions it was very successful in converging to the desired solutions. One might expect that the same happens here. And although this is true, once we include form factors and spectral densities in the bootstrap, Watson theorem provides a new natural iterative functional for our purpose. Watson theorem establishes that, when four-pion state contribution is negligible, the 2-pion form factor can be written as
$\mathcal{F}^{I}_{\ell}={}_{\text{out}}\langle\pi\pi|\mathcal{O}^{I}_{\ell}|0\rangle={}_{\text{out}}\langle\pi\pi|\pi\pi\rangle_{in}\;{}_{in}\langle\pi\pi|\mathcal{O}^{I}_{\ell}|0\rangle=S^{I}_{\ell}\mathcal{F}^{I*}_{\ell}$ (2.7)
so the phase of the form factors agrees with the phase shift of the corresponding partial wave. The procedure then replaces the phase of the previous (old) partial wave by the phase of the previous (old) form factor when computing the functional, i.e., in (2.6) we replace:
@@
Namely instead of just using the old partial wave to determine the functional, we use its modulus and replace its phase by the phase of the form factor. Such iteration will tend to converge to unitarity saturation while matching the phase of the form factor with the phase shift. From a physical point of view, the UV information incorporated into the form factors and spectral densities is conveyed to the phase shifts and partial waves. At the same time the crossing and unitarity constraints on the partial waves feed back to the form factors.
This iteratative procedure can be directly interpreted from the zero modes of the matrix. Following the form factor bootstrap [31], we introduce a $3\times 3$ positive semidefinite matrix of inner products:
@@
$G_{\text{SDP}}=\begin{array}[]{c c}&\begin{array}[]{c c c}\mbox{$|\mbox{in}\rangle$}_{P,I,\ell}&\mbox{$|\mbox{out}\rangle$}_{P,I,\ell}&{\cal O}_{P,I,\ell}\mbox{$|0\rangle$}\\ \end{array}\\ \begin{array}[]{c c c}\mbox{$\langle\mbox{in}|$}_{P^{\prime},I,\ell}\\ \mbox{$\langle\mbox{out}|$}_{P^{\prime},I,\ell}\\ \mbox{$\langle 0|$}{\cal O}^{\dagger}_{P^{\prime},I,\ell}\end{array}&\left(\begin{array}[]{c c c}1&\ S^{I}_{\ell}(s)&\ \ {\cal F}^{I}_{\ell}(s)\\ S_{\ell}^{I*}(s)&\ 1&\ \ {\cal F}^{I*}_{\ell}(s)\\ {\cal F}^{I*}_{\ell}(s)&\ {\cal F}^{I}_{\ell}(s)&\ \ \rho^{I}_{\ell}(s)\end{array}\right)\end{array}\succeq 0,\;\;\forall I,\ell,\ s>4m^{2}_{\pi},\;\;$ (2.9)
where we removed an overall factor $\delta^{4}(P-P^{\prime})$, namely the identity in the center of mass momentum. The in- and out- states are two pion states and ${\cal O}_{P,I,\ell}$ is a momentum space current. The elements of the matrix are the scattering matrix $S_{\ell}^{I}$, the form factors ${\cal F}_{\ell}^{I}$ and the spectral density $\rho_{\ell}^{I}$. Recall that for positive semidefinite Hermitian matrix $G_{\text{SDP}}$, one has
$\text{Tr}(M\cdot G_{\text{SDP}})\geq 0,\;\;\forall\;M=vv^{\dagger}$ (2.10)
@@
where $v$ is an arbitrary complex vector. When the positive semidefinite condition is saturated, this matrix has two zero eigenvalues. That means that two linear combinations of the states vanish. One simply means that the in and out states are the same (up to a phase) if there is no particle production. The other, interestingly, means that we have a way to write ${\cal O}_{I,\ell,P}\mbox{$|0\rangle$}$ as a two pion state. So, from a physical perspective, one can understand the zero modes as a way to write pQCD currents acting on vacuum in terms of pion states and vice versa. In terms of the matrix elements, this gives rise to the conditions:
$S^{I}_{\ell}=\frac{\mathcal{F}^{I}_{\ell}}{\mathcal{F}^{I*}_{\ell}},\;\;\rho^{I}_{\ell}=\mathcal{F}^{I}_{\ell}\mathcal{F}^{I*}_{\ell}$ (2.11)
where the first one is equivalent to the Watson theorem. The matrix (2.9) then becomes
$G_{\text{sat}}=\begin{pmatrix}1&\frac{\mathcal{F}_{\ell}(s)}{\mathcal{F}^{*}_{\ell}(s)}&\mathcal{F}_{\ell}(s)\\ \frac{\mathcal{F}^{*}_{\ell}(s)}{\mathcal{F}_{\ell}(s)}&1&{\mathcal{F}^{*}_{\ell}(s)}\\ {\mathcal{F}^{*}_{\ell}(s)}&{\mathcal{F}_{\ell}(s)}&{\mathcal{F}_{\ell}(s)}{\mathcal{F}^{*}_{\ell}(s)}\end{pmatrix}$ (2.12)
with two zero modes
$v_{1}=\begin{pmatrix}-\mathcal{F}_{\ell}(s)\\ 0\\ 1\end{pmatrix},\;\;v_{2}=\begin{pmatrix}-\frac{\mathcal{F}_{\ell}(s)}{\mathcal{F}^{*}_{\ell}(s)}\\ 1\\ 0\end{pmatrix}$ (2.13)
and we have
$\text{Tr}(M_{i}\cdot G_{\text{sat}})=0,\;\;M_{i}=v_{i}v_{i}^{\dagger},\;\;i=1,2$ (2.14)
i.e., the minimum attainable for (2.10). Now, given one old solution $\mathcal{F}^{(\text{old})}$, we can construct the matrices $M^{(\text{old})}=v_{i}v_{i}^{\dagger}$ following (2.13), and write down a linear functional ${\cal F}_{P}$ to minimize
$\text{min}\;\left\{{\cal F}_{P}=\int ds\sum_{I,\ell}\text{Tr}\big(M^{(\text{old})}\cdot G^{(\text{new})}\big)\right\}$ (2.15)
Repeating this procedure until convergence leads to the GTB solution, i.e., the solution we look for is a fix point of this iterative procedure. Taking $i=2$, the functional becomes precisely the Watsonian iterative functional we describe above. [^7] In addition, we can also take the other zero mode $i=1$ which we leave for future work. This procedure should be interpreted as projecting onto the null space of the positive semidefinite matrix iteratively.
@@ 2.3 Asymptotic form factors
This new improved procedure highlights the need for more precise high energy inputs. This is not a problem for the SVZ sum rules [39, 40, 41] but it is a problem for the form factors. The theoretical computation of form factors at high energy is well established [46] and proceeds similarly to the better known case of Deep Inelastic Scattering. However, there is less validation from experiments specially in the case of the pion. Qualitatively, at high energy, the pion is highly boosted but in the transverse directions it has the same size given by its radius $R_{\pi}\sim\frac{1}{f_{\pi}}$. The probability of a probing particle hitting a quark is inversely proportional to the transverse area. That is the reason that $f_{\pi}$ appears in the asymptotic formulas. More technically, the distribution amplitude is required, a low energy information. At high energy the results of Brodsky-Lepage [37, 47, 48, 3] allows us to determine form factors as
@@
2.4 SVZ sum rules
As proposed in [1, 2, 3], important information about the gauge theory is incorporated into the bootstrap by means of the SVZ sum rules [39, 40, 41]. In this paper, as compared to the previous ones we use a slightly improved version of the finite energy sum rules as is explained in detail in Appendix C where we also quote 3-loops results taken from the available literature on perturbative QCD.
@@ 6 Thermodynamics of a dilute gas of pions, second virial coefficients and density of pion pairs
6.1 Pairs of pions with opposite and same sign of electric charge
A quantity that is of interest for pions produced in high energy nuclear collisions [^10] We are grateful to Fuqiang Wang for a discussion on this topic. [67], when the quark gluon plasma cools down, is the difference in the pairs of pions with opposite (electric) charge $N_{\mbox{os}}$ minus the number of pairs with the same sign $N_{\mbox{ss}}$. If we consider that quantity in a gas at equilibrium, as we do here, then we define
$N_{\mbox{os}}-N_{\mbox{ss}}=2\langle\langle N_{+}N_{-}\rangle\rangle-\langle\langle N_{+}N_{+}\rangle\rangle-\langle\langle N_{-}N_{-}\rangle\rangle$ (6.27)
where the brackets $\langle\langle\rangle\rangle$ indicate thermal average. For a dilute gas of free pions this difference vanishes since e.g. $\langle\langle N_{+}N_{-}\rangle\rangle=\langle\langle N_{+}\rangle\rangle\langle\langle N_{-}\rangle\rangle$ and $\langle\langle N_{+}\rangle\rangle=\langle\langle N_{-}\rangle\rangle$. In an interacting gas, this is no longer true, the eigenstates of the Hamiltonian that enter the partition function are not eigenstates of $N_{\pm}$. Suppose we want to compute the number of pairs with opposite sign
$N_{OS}=\langle\langle N_{+}N_{-}\rangle\rangle=\frac{1}{\Xi}\sum_{\nu}N_{+}N_{-}e^{-\beta E_{\nu}+\beta\mu N_{\nu}}$ (6.28)
Expanding in powers of the fugacity it is clear we need at least two particles to have a pair so we get, at order $z^{2}$:
$N_{OS}-N_{OS}^{free}$ $=$ $z^{2}\sum_{\nu,N_{\nu}=2}N_{+}N_{-}e^{-\beta E_{\nu}}$
$=$ $Vz^{2}\int dM\frac{M^{2}}{2\pi^{2}\beta}\,K_{2}(\beta M)\sum_{\ell,I,I_{z}}(2\ell+1)\mbox{$\langle II_{z}|$}N_{+}N_{-}\mbox{$|II_{z}\rangle$}\frac{1}{\pi}\frac{\partial\delta^{I}_{\ell}}{\partial M}$
In the sum over $I_{z}$ only $I_{z}=0$ survives. Now we have
[…]
whereas for same sign we have $I_{z}=2$ and then $I=2$. Putting everything together we find
$N_{OS}-N_{SS}$ $=$ $Vz^{2}\int_{2m_{\pi}}^{\infty}dM\frac{M^{2}}{2\pi^{2}\beta}\,K_{2}(\beta M)\sum_{\ell}{}^{\prime}(2\ell+1)\frac{1}{\pi}\frac{\partial}{\partial M}\left(\frac{1}{3}\delta^{0}_{\ell}+\frac{1}{2}\delta^{1}_{\ell}-\frac{5}{6}\delta^{2}_{\ell}\right)$
where $\sum{}^{\prime}$ indicates that the sum is over even or odd $\ell$ for even or odd isospin. Still using the distribution in $M$ ($M^{2}=(p_{1}+p_{2})^{2}$) but introducing $s=M^{2}$ we get:
$N_{OS}-N_{SS}=$ $Vz^{2}\int_{2m_{\pi}}^{\infty}dM\;n_{os-ss}(M)$ (6.31)
$=$ $Vz^{2}\int_{2m_{\pi}}^{\infty}dM\frac{s^{\frac{3}{2}}}{\pi^{3}\beta}K_{2}(\beta\sqrt{s})\frac{\partial}{\partial s}\left(\frac{1}{3}\delta_{S0}+\frac{3}{2}\delta_{P1}-\frac{5}{6}\delta_{S2}+\frac{5}{3}\delta_{D0}\right)$
The results are depicted in fig. 15 for the case of zero chemical potential $z=1$. The $\rho$ resonance peak coming from $\frac{\partial\delta_{1}^{1}}{\partial M}$ is evident. It can be interpreted as an increase in the density of states near the $\rho$ mass or as $\rho$ mesons being created and destroyed leading in both cases to an extra density of pairs of opposite sign.